OpenStudy (anonymous):

Multiply: 6y√7 x 2√y Help Plzzzzzzzzzzz.!

OpenStudy (anonymous):

multiply all the number

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

\[(6*2) \sqrt{7} y \sqrt{y}\] \[12 \sqrt{7} y \sqrt{y}\]

OpenStudy (anonymous):

Thanks.!